# Identifying the impact of local connectivity patterns on dynamics in excitatory-inhibitory networks

### v2

@@ Acknowledgements.
YS and SO were supported by the Eranet-Neuron project IMBALANCE and the program “Ecoles Universitaires de Recherche” launched by the French Government and implemented by the ANR, with the reference ANR-17-EURE-0017. YS was supported by the National Natural Science Foundation of China (grant no. 32400936), the Fundamental Research Funds for the Central Universities (grant no. 2233100030), the NSF AccelNet IN-BIC program (grant No. OISE-2019976). DD was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of the SPP 2205 – 533396241. We additionally acknowledge partial support of NIH grant R01 1RF1DA055669. We thank K.D. Miller for comments on the manuscript.
@@ Appendix A Other second-order motifs
$\frac{\left[\left(\mathbf{Z}^{2}\right)_{ij}\left(\mathbf{Z}^{2}\right)_{jk}\right]}{\left[\mathbf{Z}^{2}\right]_{ij}\left[\mathbf{Z}^{2}\right]_{jk}}$ $=1+\frac{N\tau^{div}\tau^{con}\sigma_{z}^{4}+N\tau^{c}\tau^{r}\sigma_{z}^{4}}{(N^{2}\tau^{c})^{2}\sigma_{z}^{4}}$ (58)
@@ B.3 Gaussian networks with heterogeneous variance and correlations
Appendix C Network constructions
@@ C.2 Sparse networks
Appendix D Supplementary figures.
@@
Figure D.1: Eigenvalues of networks featuring reciprocal motifs. In subplots (a,b), results are presented for networks with Gaussian-distributed synaptic couplings. Subplots (a) and (b) show how the unique eigenvalue outlier $\lambda_{1}$ and the radius of the eigenvalue bulk $R_{bulk}$ change as the statistics of the reciprocal motifs $\tau^{r}$ increase. The red error bars with asterisks represent the numerically obtained $\lambda_{1}$ from 30 network realizations, the red lines represent the theoretical predictions obtained using Eqs.(34) and (36). Gray areas illustrate the area of eigenvalue bulks obtained numerically, with gray dash-lines representing the theoretical predictions [30]. The black dash-lines indicate the unperturbed eigenvalue $\lambda_{0}$ of the mean connectivity. In subplot (a), the network parameters are the same as those used in Fig 2: $\sigma=0.2$ and $\tau^{r}$ ranges from 00 to $0.1$. In subplot (b), the networks have parameters $\sigma=0.5$, and $\tau^{r}$ ranges from 00 to $0.8$. Subplots (c) exhibit results for sparse networks. Figure descriptions are consistent with those in subplots (a,b).
@@
Figure D.2: Mean response $\chi_{E}$ of the excitatory neuron population in response to the time-invariant uniform input $I^{ext}$ in networks with only reciprocal motifs. Subplots (a) corresponds to Gaussian networks, while subplots (b) corresponds to sparse networks. Network parameters remain consistent with those provided in Figs 6 and 9.
@@
Figure D.3: Population-averaged mean of entries on connectivity vectors for sparse networks featuring chain motifs. Figure descriptions are identical to those in Fig. 8, except that the network parameters are consistent with those used in Fig. 10 (see TABLE 2).
@@
Figure D.4: Impact of chain motifs on the mean response $\chi_{pq}$ of population $p$ to uniform inputs to population $q$, in sparse excitatory-inhibitory networks. Figure descriptions are identical to those in Fig. 10, except that the network parameters are consistent with those used in Figs 89 (see TABLE 2).

### v3

@@ Acknowledgements.
YS and SO were supported by the Eranet-Neuron project IMBALANCE and the program “Ecoles Universitaires de Recherche” launched by the French Government and implemented by the ANR, with the reference ANR-17-EURE-0017. YS was supported by the National Natural Science Foundation of China (grant no. 32400936), the Fundamental Research Funds for the Central Universities (grant no. 2233100030), the NSF AccelNet IN-BIC program (grant No. OISE-2019976). DD was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of the SPP 2205 – 533396241. We additionally acknowledge partial support of NIH grant R01 1RF1DA055669. We thank K.D. Miller for feedback on the manuscript and the suggestion to look for an equivalent effective matrix.
@@ Appendix A Other second-order motifs
$\frac{\left[\left(\mathbf{Z}^{2}\right)_{ij}\left(\mathbf{Z}^{2}\right)_{jk}\right]}{\left[\mathbf{Z}^{2}\right]_{ij}\left[\mathbf{Z}^{2}\right]_{jk}}$ $=1+\frac{N\tau^{div}\tau^{con}\sigma_{z}^{4}+N\tau^{c}\tau^{r}\sigma_{z}^{4}}{(N\tau^{c})^{2}\sigma_{z}^{4}}$ (64)
@@ B.3 Gaussian networks with heterogeneous variance and correlations
Appendix C Network construction
@@ C.2 Sparse networks
Appendix D Effective connectivity matrix for the response function
In this appendix, we show that the response function obtained from the effective connectivity $\mathbf{J}^{eff}=\mathbf{J^{0}}+[\mathbf{Z}^{2}]$ is identical at leading order in $N$ to the mean response $[\bm{\chi}]$ obtained by averaging Eq. (10) over realizations of the random connectivity.
We begin with Eq. (10), and use the decomposition $\mathbf{J}=\mathbf{J^{0}}+\mathbf{Z}$ (Eq. (4)). The mean response function $[\bm{\chi}]$ is expressed as
$[\bm{\chi}]=\left[(\mathbf{1}-\mathbf{Z}-\frac{1}{N}\mathbf{M}_{0}\mathbf{N}_{0}^{\intercal})^{-1}\right],$ (98)
where the mean connectivity $\mathbf{J^{0}}=\frac{1}{N}\mathbf{M}_{0}\mathbf{N}_{0}^{\intercal}$ has a block structure with constant elements, and is also low-rank. Using the Woodbury matrix identity and expanding the matrix inverse as a power series [35], we have
$\left[\bm{\chi}\right]$ $=\left[(\mathbf{1}-\mathbf{Z}-\frac{1}{N}\mathbf{M}_{0}\mathbf{N}_{0}^{\intercal})^{-1}\right]$ (99)
$=\left[(\mathbf{1}-\mathbf{Z})^{-1}\right]-\bigg{[}(\mathbf{1}-\mathbf{Z})^{-1}\left(-\frac{\mathbf{M}_{0}}{N}\right)\left(\mathbf{1}-{\mathbf{N}_{0}^{\intercal}(\mathbf{1}-\mathbf{Z})^{-1}\frac{\mathbf{M}_{0}}{N}}\right)^{-1}\mathbf{N}_{0}^{\intercal}(\mathbf{1}-\mathbf{Z})^{-1}\bigg{]}$
$\approx\left[(\mathbf{1}-\mathbf{Z})^{-1}\right]+\bigg{[}(\mathbf{1}-\mathbf{Z})^{-1}\frac{\mathbf{M}_{0}}{N}\mathbf{N}_{0}^{\intercal}(\mathbf{1}-\mathbf{Z})^{-1}\bigg{]}$
$\approx\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]\right)^{-1}+\bigg{[}\sum_{k,l=0}^{\infty}\mathbf{Z}^{k}{\frac{\mathbf{M}_{0}\mathbf{N}_{0}^{\intercal}}{N}}\mathbf{Z}^{l}\bigg{]}$
$=\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]\right)^{-1}+\sum_{k,l=0}^{\infty}\left[\mathbf{Z}^{2}\right]^{k}\frac{\mathbf{M}_{0}\mathbf{N}_{0}^{\intercal}}{N}\left[\mathbf{Z}^{2}\right]^{l}.$
From the second to the third line, we keep only the first order in $\frac{1}{N}$. From the second-to-last to the last line, we note that the low-rank mean connectivity has a block-like structure with deterministic elements, whereas $\mathbf{Z}$ consists of zero-mean random elements. So $\mathbf{Z}$ is independent of the low-rank mean connectivity, and only terms of the form $[\mathbf{Z}^{2}]^{k}$ remain in the large network limit.
We next consider the response function $\bm{\chi}^{eff}$ generated by $\mathbf{J}^{eff}=\mathbf{J^{0}}+[\mathbf{Z}^{2}]$ and perform the same series of manipulations:
$\bm{\chi}^{eff}$ $=\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]-\frac{1}{N}\mathbf{M}_{0}\mathbf{N}_{0}^{\intercal}\right)^{-1}$ (100)
$=\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]\right)^{-1}-\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]\right)^{-1}\left(-\frac{\mathbf{M}_{0}}{N}\right)\left(\mathbf{1}-{\mathbf{N}_{0}^{\intercal}\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]\right)^{-1}\frac{\mathbf{M}_{0}}{N}}\right)^{-1}\mathbf{N}_{0}^{\intercal}\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]\right)^{-1}$
$\approx\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]\right)^{-1}+\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]\right)^{-1}\frac{\mathbf{M}_{0}}{N}\mathbf{N}_{0}^{\intercal}\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]\right)^{-1}$
$=\left(\mathbf{1}-\left[\mathbf{Z}^{2}\right]\right)^{-1}+\sum_{k,l=0}^{\infty}\left[\mathbf{Z}^{2}\right]^{k}\frac{\mathbf{M}_{0}\mathbf{N}_{0}^{\intercal}}{N}\left[\mathbf{Z}^{2}\right]^{l}.$
Therefore, at leading order in $N$, we find that the averaged response function $[\bm{\chi}]$ obtained using Eq. (99) is identical to the response function $\bm{\chi}^{eff}$ derived from the effective connectivity $\mathbf{J}^{eff}=\mathbf{J^{0}}+[\mathbf{Z}^{2}]$.
We test the accuracy of $\bm{\chi}^{eff}$ in both Gaussian (Figs. 6 and 7) and sparse networks (Figs. 9, 10 and 11 on the right) in the main text. The results show that the response function derived using $\mathbf{J}^{eff}$ is in good agreement with the response function derived using either numerical simulations or those obtained using the low-rank framework.
Examining the eigenvalues of the effective connectivity matrix $\mathbf{J}^{eff}$, we find that they differ from the outliers of the full connectivity matrix $\mathbf{J}$. For instance, for fully connected networks with uniform variance (Sec. III.2) $\mathbf{J}^{eff}$ has a unit rank structure, and therefore only one non-zero eigenvalue, while $\mathbf{J}$ has two outliers as $\tau^{c}$ is increased (Fig. 2). The single unique eigenvalue is given by
$\lambda_{eff}=\lambda_{0}+N\sigma^{2}\tau^{c},$ (101)
and therefore increases linearly with $\tau^{c}$, while the positive outlier of $\mathbf{J}$ increases as $\sqrt{\tau^{c}}$ (Eqs. (34-35)).
Despite this difference, we find that $\mathbf{J}^{eff}$ predicts correctly network instability. Indeed for the full connectivity matrix $\mathbf{J}$, the positive outlier crosses unity when
$\frac{\lambda_{0}+\sqrt{\lambda_{0}^{2}+4(N-1)\sigma^{2}\tau^{c}}}{2}=1,$ (102)
here we use $N$ instead of $N-1$ (Eqs. (34), (35)) in the large network limit. Rearranging terms, this is directly equivalent to $\lambda_{eff}=1$. Therefore, the effective connectivity matrix $\mathbf{J}^{eff}$ accurately predicts the instability of the network dynamics.
Appendix E Supplementary figures.
@@ Appendix D Effective connectivity matrix for the response function
Figure E.1: Eigenvalues of networks featuring divergent, convergent and reciprocal motifs. In subplots (a-d), results are presented for networks with Gaussian-distributed synaptic couplings. Subplots (a), (b), (c) and (d) show how the unique eigenvalue outlier $\lambda_{1}$ and the radius of the eigenvalue bulk $R_{bulk}$ change as the statistics of the divergent $\tau^{div}$, convergent $\tau^{con}$ and reciprocal motifs $\tau^{r}$ increase. The red error bars with asterisks represent the numerically obtained $\lambda_{1}$ from 30 network realizations, the red lines represent the theoretical predictions. Gray areas illustrate the area of eigenvalue bulks obtained numerically, with gray dashed lines representing the theoretical predictions [30]. The black dashed lines indicate the unperturbed eigenvalue $\lambda_{0}$ of the mean connectivity. In subplot (a-c), the network parameters are the same as those used in Fig 2: $\sigma=0.2$ and $\tau^{div/con/r}$ ranges from 00 to $0.1$. In subplot (d), the networks have parameters $\sigma=0.5$, and $\tau^{r}$ ranges from 00 to $0.8$. Subplot (e) exhibits results for sparse networks. Figure descriptions are consistent with those in subplots (a-d).
@@
Figure E.2: Mean response $\chi_{E}$ of the excitatory neuron population in response to the time-invariant uniform input $I^{ext}$ in networks with only reciprocal motifs. Subplots (a) corresponds to Gaussian networks, while subplots (b) corresponds to sparse networks. Network parameters remain consistent with those provided in Figs 6 and 9.
@@
Figure E.3: Population-averaged mean of entries on connectivity vectors for sparse networks featuring chain motifs. Figure descriptions are identical to those in Fig. 8, except that the network parameters are consistent with those used in Fig. 10 (see TABLE 2).
@@
Figure E.4: Impact of chain motifs on the mean response $\chi_{pq}$ of population $p$ to uniform inputs to population $q$, in sparse excitatory-inhibitory networks. Figure descriptions are identical to those in Fig. 10, except that the network parameters are consistent with those used in Figs 89 (see TABLE 2).
Figure E.5: Distinct effects of EE-type (panel (a), $\tau^{c}_{EE}>0$), EI-type (panel (b), $\tau^{c}_{EI}<0$), IE-type (panel (c), $\tau^{c}_{IE}<0$), and II-type (panel (d), $\tau^{c}_{II}$) chain motifs on eigenvalues (1.), inhibitory-associated low-rank structures (2.), and response functions $\chi_{E}$ and $\chi_{II}$ (3.). Results are obtained using the Gaussian approximation for sparse networks; except for the varying chain-motif strength, network parameters are identical to those in Figs.10 and11.
Figure E.6: Distinct effects of EI-type (panel (a), $\tau^{c}_{EI}>0$) and IE-type (panel (b), $\tau^{c}_{IE}>0$) chain motifs on eigenvalues (1.), inhibitory-associated low-rank structures (2.), and response functions $\chi_{E}$ and $\chi_{II}$ (3.). Other figure descriptions remain consistent with those in Fig. E.5.